With an irradiance of 0.001W/cm2 incident on a detector of area 0.5 cm 2 and with a bandwidth of 2 Hz, the ratio of the noise voltage to the signal voltage is 10. What is the noise equivalent power of the detector?
Answer #1

Solution:

Given:

Irradiance, H=0.001 W/cm2

Area, A=0.5 cm2

Bandwidth= 2Hz

Ratio of the noise voltage to the signal voltage is 10

We know that Noise to Equivalent Power ratio is given by,

NEP= \frac{HAV_{_{N}}}{V_{S}(\Delta f)^{\frac{1}{2}}}

Where,

H= Irradiance

A= Area

VS= Root mean square signal voltage

VN = Root mean square noise voltage within the measurement bandwidth ∆f

Using the above formula we get,

NEP= \frac{0.001*0.5*10}{(2)^{\frac{1}{2}}}

NEP= 0.003535534 W/Hz^{\frac{1}{2}}

=3.5\times 10^{-3}W/Hz^{\frac{1}{2}}